OpenStudy (anonymous):

Identify the excluded value for each function: 34. f(x)=5/x 35. y=3/x+4

OpenStudy (campbell_st):

a zero in the denominator of a fraction is undefined. So the question is asking you to identify what number can't be put into the denominator of each fraction. as an example \[\frac{5}{x - 10}\] is x = 10 the denominator becomes 10 - 10 = 0 so x = 10 is an excluded value. Hope this helps